Solve the inequality x+14/x-8>=1
Answer by graphmatics(170)

x+14/x-8>=1
Let x be greater then 0 then:
x^2+14-8x>=x
x^2-9x+14>=0
(x-2)(x-7)>=0 So
x-2>=0 So x>=2
x-7>=0 So x>=7 We must select x>=7 or x=2. There is a similar set of inequalities when we let x be less than 0